Maximize U(x,y)=xy+x+2y subjected to Px=,Py=5 and M=51
1
Expert's answer
2021-08-02T11:04:30-0400

Solution:

Utility function U(x,y) = xy + x + 2y

The utility-maximizing condition: "\frac{MUx}{MUy} = \frac{Px}{Py}"


First, derive MRS:

MRS = "\frac{MUx}{MUy}"


MUx = "\frac{\partial U} {\partial X} = y + 2y = 3y"


MUy = "\frac{\partial U} {\partial Y} = x + x = 2x"


MRS = "\frac{MUx}{MUy} = \frac{3y}{2x}"


Set MRS equal to "\frac{Px}{Py}" to derive the utility-maximizing bundle:

Px = 2

Py = 5

"\frac{3y}{2x} = \frac{2}{5}"


y = "\frac{4x}{15}"


Plug Y into the budget constraint to derive X:

Budget constraint: M = PxX + PyY

51 = 2X + 5Y

51 = 2X + 5("\frac{4x}{15})"


51 = 2X + "\frac{20x}{15}"


51 = 2X + "\frac{4x}{3}"

Multiple all sides by 3

153 = 6X + 4X

153 = 10X

X = 15.3

Plug this into Y equation:

Y = "\frac{4x}{15} = \frac{(4\times 15.3)}{15} = \frac{61.2}{15} = 4.08"

Y = 4.08


Utility maximizing bundle (Ux,y) = (15.3, 4.08)